Grade 12Mechanics

a ball is dropped from the top of a building.the ball takes 0.5 sec to fall past the 3m height of a window some distance from the top of building.if the speed of the ball at the top and at the bottom of the window are v and u then
a.v+u=12m/s b.v-u=4.9m/s
c.vu=1m/s d.v/u=1m/s

Gaurav Talekar
10 Years agoGrade 12

1 Answer

Madhavan
8 Years ago
If we take g=10 m/s2,
Given that,t=0.5seconds,s=3m,
If we apply it in the equation of
S=(v+u)÷2×t,we get,3=(v+u)÷2×10.
Thus we get v+u=12m/s